Find the first negative term of an AP 129,120,111,102....and so on.

Answer by rothauserc(4718)

the nth term of an arithmetic progression is
Xn = a1 + d(n-1)
for our problem a1=129, d=-9
Xn = 129 - 9(n-1) < 0 =
-9(n-1) < -129 =
-9n+9 < -129 =
-9n < -138
n > 15.3, therefore
the delta must be 9*16
X17 = 129 - 9(17-1)
the 17th term of the arithmetic progression is the first term to be negative